Suppose you structured a bond deal for Kanye West that paid him $20 million upfront, with future royalties and streaming revenues from his past albums going towards payments to bondholders. Each bond had a face value of $1,000 and a coupon rate of 7.7% with semi-annual coupons. If the bonds have 9 years remaining until maturity and the current yield to maturity is 9.7%, what price is each bond trading at right now? Round to the nearest cent.
Answer #1

Now, for this question,

M = $1000, n = 9 * 2 = 18 semi-annual periods, C = 7.7% * $1000/2 = $38.5 (semi-annually), i = 9.7%/2 = 4.85% (semi-annually)

P = $455.37 + $$426.35

P = $881.7 ----> Answer